OpenStudy (anonymous):

Please help If eight times a number is added to 12, the result is three less than three times the number. What is the number?

OpenStudy (anonymous):

so do 8x+12=3x-3

OpenStudy (anonymous):

let the number be x given 8x + 12 = 3x - 3 8x - 3x = -3 - 12 5x = -15 x = -3

OpenStudy (anonymous):

do you get that @natth88 ?